OpenStudy (anonymous):

pls help. Choose the property which best suits each step. (1/5) [(n+5) + -n] Given =(1/5) [(5 + n) + -n] A =(1/5) [ 5 + (n + -n)] B =(1/5) [5+0] C =(1/5)(5) D =1 E 1) Additive Inverse ___ 2) Commutative ___ 3) Associative ___ 4) Multiplicative Inverse ___ 5) Additive Identity ___ ive tried looking definitions up and I dont seem to understand. thanks in advanced..

OpenStudy (anonymous):

Additive Inverse: a+(-a)=0 Commutative: a+b = b+a Associative: a+(b+c) = (a+b)+c Multiplicative Inverse: x*(1/x) = 1 Additive Identity: a+0=a 1) C 2) A 3) B 4) E 5) D

OpenStudy (anonymous):

@JessCALI Knock knock are you there?

OpenStudy (anonymous):

Thank you so much..